OpenStudy (anonymous):

A motorboat takes 5 hours to travel 450 mi going upstream. The return trip takes 3 hours going downstream. What is the rate of the boat in still water and what is the rate of the current?

OpenStudy (ganpat):

x = speed of boat y = speed of current U = rate upstream V = rate Downstream.. U = (x - y) = 450/ 5 = 90.. V = (x + y) = 450/ 3 = 150.. So rate of boat is still water = (U + V)/ 2 = 240 /2 = 120 mph.. rate of current= 120 - y = 90 = 30 mph..

OpenStudy (anonymous):

I don't understand what you're saying

OpenStudy (ganpat):

i used the basic formula , Speed = distance / time.. Now we do not have only speed of boat.. But the boat is facing currents i.e. water speed. So when it is Upstream = x- y ..... and for downstream = x + y getting ??